OpenStudy (anonymous):

Find the midpoint of the line segment with the given endpoints. (-7/3, 3/4), (5/3, -9/4)

OpenStudy (wolf1728):

Averaging the x values (-2.3333 + 1.6666) / 2 x = -.3333333 = -1/3 y = (.75 -2.25) / 2 y = -.75 = -3/4

OpenStudy (anonymous):

Let the mid point be M(x.y)-----X=(-7/3+5/3)/2=(-2/3)/2=-1/3 Y=(-9/4+3/4)/2=(-3/2)/2=-3/4 Therefore M(x,y)=(-1/3,-3/4)